OpenStudy (anonymous):

A survey showed that 26% of the people surveyed preferred a certain brand of breakfast cereal. There were 200 people surveyed. How many people surveyed preferred the brand of breakfast cereal? A. 13 B. 26 C. 52 D. 126

OpenStudy (anonymous):

@KyanTheDoodle @KendrickLamar2014

OpenStudy (anonymous):

@shifuyanli

OpenStudy (anonymous):

what do you think @TheAsker2002

OpenStudy (anonymous):

b

OpenStudy (anonymous):

nope your wrong

OpenStudy (anonymous):

can you help...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Tell mw how to do it please

OpenStudy (anonymous):

me*

OpenStudy (anonymous):

just multiply the number if 26 by 2.

OpenStudy (anonymous):

52

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

if it was out of 100 people it would be 26 for 200 just do 26 * 2

OpenStudy (kendricklamar2014):

13/200 = 6.5% 26/200 = 13% 52/200 = 26%

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

multiply 200 x 0.26